OpenStudy (anonymous):

find the system of equations -x+y-z=0 x+2y+z=9 y+z=5

OpenStudy (agent0smith):

-x+y-z=0 x+2y+z=9 y+z=5 You can add together the first two equations, to eliminate y and z. After adding the first two, you have: -x+y-z=0 3y=9 (our new second equation after adding the first and second) y+z=5 Now notice you can just solve the last two equations: 3y=9 y+z=5

OpenStudy (anonymous):

so y =3

OpenStudy (anonymous):

and z=2

OpenStudy (anonymous):

x=-1.?

OpenStudy (agent0smith):

Yes, yes, no. x will be a whole number. Use -x+y-z=0 to find x, it's the easiest way.

OpenStudy (agent0smith):

Oh, you said x=1, yes that's right.